# Survival Modeling from Whole Slide Images via Patch-Level Graph Clustering and Mixture Density Experts

### v4

@@ 1 Introduction
Although MIL has proven effective for cancer grading and subtyping, survival prediction further requires modeling long-range dependencies and subtle morphological cues across distant tissue regions. Traditional methods like the Cox proportional hazards model fail to capture such complex, non-linear relationships, while deep learning models often lack interpretability and spatial awareness.
@@ 3.2 Graph-Regularized Patch Clustering (GRPC)
Using the similarity matrix $S$, a $k$-nearest-neighbor (k-NN) graph is constructed, and GPU-accelerated K-Means guided by the graph topology clusters patches into $K$ groups $L_{1},\dots,L_{K}$}. Each cluster reflects a specific tissue microenvironment—like tumor epithelium, stroma, or necrosis—providing a biologically meaningful and spatially consistent structure within the slide.
@@ 3.3 Hierarchical Feature Aggregation (HFA)
Intra-cluster features are concatenated as $\widetilde{P}=\mathrm{Concat}(L^{\prime}_{k})_{k=1}^{K}$. Averaging $R^{\prime}$ across clusters produces a broadcast descriptor $R^{\prime}_{\mathrm{exp}}$, expanded to match $\widetilde{P}$ and added residually to yield $\widehat{P}$. Finally, $\widehat{P}$ is concatenated with unselected irrelavant patch features to form $\mathcal{P}_{\text{final}}$.
@@ 3.4 Expert-Guided Mixture Density Modeling
An Expert-Guided Mixture Density Modeling block (Figure 1b) is utilized to estimate continuous survival probabilities by modeling the complete survival time distribution. Based on the mixture-of-experts paradigm [bishop2006pattern] and SurvMDN [han2022survivalmixturedensitynetworks], it employs multiple experts to construct Gaussian mixtures that represent diverse survival patterns.
@@ 3.5 Training Objective and Regularization
To promote interpretability and specialization, a gating entropy loss is introduced ($\mathcal{L}_{\text{ent}}$).
@@
Table 1: Left: TCGA results (TDC, higher is better). Right: Ablations on $q$ in QPF, GRPC, and HFA. Best TDC values in bold.
@@ 4.2 Training Configuration
The quantile of pacth filtering module is fixed at $q=0.25$, selecting the top $75\%$ of patches by their logits. In the k-NN graph, each patch links to 10 neighbors, which are grouped into $C$ clusters and processed through an 8-head hierarchical attention module. The Expert-Guided Mixture Density block employs five experts, each fitting a Gaussian Mixture Model (GMM) with $K=100$ components. Training runs for 20 epochs using Adam (lr = $2\times 10^{-4}$, weight decay = $1\times 10^{-3}$, dropout = 0.1) with a batch size of 1. Results are reported as mean $\pm$ standard deviation of the Time-Dependent Concordance Index (TDC) [han2022survivalmixturedensitynetworks] metric across 5-fold cross-validation, using an NVIDIA A6000 GPU.
@@ 5.2 Ablation Studies and Interpretability
Ablation experiments were conducted to assess the impact of Quantile-Gated Patch Filtering, Graph-Regularized Patch Clustering, and Hierarchical Feature Aggregation (Table 1(b)). Varying the quantile showed that q=0.25q{=}0.25 offers the best trade-off, preserving prognostic regions while removing noise. Eliminating any module reduces TDC, confirming their complementary roles in denoising and spatial reasoning. For interpretability, patients were stratified into high- and low-risk groups using predicted survival scores, and Kaplan–Meier curves were plotted (Fig. 2). The curves show clear survival separation (log-rank p=0.049p{=}0.049 for LUAD, p=0.047p{=}0.047 for KIRC, p=0.089p{=}0.089 for BRCA), demonstrating that the expert-guided design captures prognostic morphology. Dynamic patch selection and graph-aware attention further enhance high-risk region localization and interpretability.

### v6

@@ 1 Introduction
Although MIL has proven effective for cancer grading and subtyping, survival prediction further requires modeling long-range dependencies and subtle morphological cues across distant tissue regions. Traditional methods like the Cox proportional hazards model fail to capture such complex, non-linear relationships, while previous deep learning models often lack interpretability and spatial awareness.
@@ 3.2 Graph-Regularized Patch Clustering (GRPC)
Using the similarity matrix $S$, a $k$-nearest-neighbor (k-NN) graph is constructed, and GPU-accelerated K-Means guided by the graph topology clusters patches into $G$ groups $L_{1},\dots,L_{G}$}. Each cluster reflects a specific tissue microenvironment—like tumor epithelium, stroma, or necrosis—providing a biologically meaningful and spatially consistent structure within the slide.
@@ 3.3 Hierarchical Feature Aggregation (HFA)
Intra-cluster features are concatenated as $\widetilde{P}=\mathrm{Concat}(L^{\prime}_{g})_{g=1}^{G}$. Averaging $R^{\prime}$ across clusters produces a broadcast descriptor $R^{\prime}_{\mathrm{exp}}$, expanded to match $\widetilde{P}$ and added residually to yield $\widehat{P}$. Finally, $\widehat{P}$ is concatenated with unselected irrelavant patch features to form $\mathcal{P}_{\text{final}}$.
@@ 3.4 Expert-Guided Mixture Density Modeling
The Expert-Guided Mixture Density Modeling block (Figure 1b) estimates continuous survival probabilities by modeling survival times, with each expert focusing on distinct WSI features $\mathbf{z}_{\text{WSI}}$. Following the mixture-of-experts paradigm [bishop2006pattern] and SurvMDN [han2022survivalmixturedensitynetworks], it constructs Gaussian mixtures to represent diverse survival patterns.
@@ 3.5 Training Objective and Regularization
To promote interpretability and specialization, a gating entropy loss ($\mathcal{L}_{\text{ent}}$) is introduced for diverse expert usage.
@@
Table 1: Left: TCGA results (TDC, higher is better). Right: Ablations on $q$ in QPF, GRPC, HFA and $e$ in EGMDM. Best TDC values in bold.
@@ 4.2 Training Configuration
The quantile of patch filtering module is fixed at qq=0.25, selecting the top $75\%$ of patches by their logits. The k-NN graph built on similarity matrix $S$, connects each node to 10 neighbors, clustering patch indices into $G$ groups of 64 each and those corresponding patches are processed by an 8-head MHSA in the HFA module. The Expert-Guided Mixture Density block employs five experts, each modeling a K=100K\!=\!100 Gaussian mixture. Training runs for 20 epochs with Adam (lr $2\times 10^{-4}$, weight decay $1\times 10^{-3}$), batch size 1. Performance is reported as mean $\pm$ std of the Time-Dependent Concordance Index (TDC) [han2022survivalmixturedensitynetworks] over 5-fold cross-validation on an NVIDIA A6000 GPU.
@@ 5.2 Ablation Studies and Interpretability
Ablation experiments evaluated the effects of Quantile-Gated Patch Filtering, Graph-Regularized Patch Clustering, Hierarchical Feature Aggregation, and Expert-Guided Mixture Density Modeling (Table 1(b)). A quantile of q=0.25q{=}0.25 offered the best trade-off, retaining prognostic regions while suppressing noise. Removing any module reduced TDC, confirming their complementary roles in denoising and spatial reasoning. Single-expert models underperformed, whereas five experts improved results, highlighting the value of multi-expert modeling. For interpretability, patients were stratified into high- and low-risk groups using predicted survival scores. Kaplan–Meier curves (Fig. 2) showed clear separation (log-rank p=0.049p{=}0.049 LUAD, p=0.047p{=}0.047 KIRC, p=0.089p{=}0.089 BRCA), validating prognostic relevance and regional localization.